from a uniform circular sheet disc of radius 2cm(its centre of mass is at O), a circular portion of radius 1cm is removed such that shift in centre of mass is maximum. the disc is now rotated about O perpendicular to the plane through α, then the magnitude of displacement of new centre of mass is 1/√3 cm, then α is
Ans 120°
1
Expert's answer
2020-09-21T08:29:57-0400

As "m_1\sdot b=m_2\sdot r" where "b=" OC, hence "b=\frac{m_2r}{m_1}=\frac{\pi\sdot r^2r}{\pi(R^2-r^2)}=\frac{r^3}{R^2-r^2}=1/3" where "R=2cm" , OO'"=" "r=1cm" , then the magnitude of displacement CC' is "CC'=\sqrt{b^2+b^2-2b^2\cos\alpha}=(2/3)\sin\alpha/2" . As "CC'=1/\sqrt3" hence "\sin\alpha/2=\sqrt3/2" hence "\alpha=120^0"